Circular No. 7231

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V348 SAGITTARII A. U. Landolt, Louisiana State University; and A. K. Uomoto, Johns

Hopkins University, observing with a CCD at the Las Campanas Observatory 2.5-m du Pont telescope, find the following preliminary V magnitudes for this unusual object (cf. IAUC 5640): “July 31.18 UT, 18.10; Aug. 1.15, 18.18; 2.18, 18.05; 3.19, 17.63; and 4.19, 17.66. The data indicate that the star may have begun a recovery from its most recent minimum.”

COMET P/1999 P1 (MACHHOLZ 2) R. H. McNaught reports his recovery of comet P/1994 P1 (= 1994o

= 1994 XXVI) on CCD images obtained with the 1.0-m f/8 reflector at Siding Spring. The object is of stellar appearance. The indicated correction to the prediction by B. G. Marsden on MPC 27082 (for component A) is ΔT = +0.8 day. Seeing was good on Aug. 4, and there was no sign of any other components within ΔT = ±1.5 days.

1999 UT  α_{2000}  δ_{2000}  m_{2}
Aug. 3.55335  16^{h}49^{m}18.^{s}57  −11^{o}11′12′′.3  20.8
3.55958  16 49 18.26  −11 11 11.7  20.3
3.57818  16 49 17.25  −11 11 07.9  20.4
4.42617  16 48 36.00  −11 08 49.8
4.42864  16 48 35.88  −11 08 49.4
4.43115  16 48 35.76  −11 08 49.1
4.43396  16 48 35.61  −11 08 48.6

Further orbital computations by the undersigned confirm that if the above observations are of the same object that was observed at Siding Spring on 1995 Mar. 29 and 30 (MPC 25097), this is indeed component A. However, attempts to link all the observations (back to 1994 Aug. 15), even using the nongravitational parameters A_{1} and A_{2}, have not been satisfactory. But the following gravitational solution gives an acceptable fit to 67 observations back to 1994 Oct. 2 (mean residual 0′′.9; earlier residuals increasing to 20′′):

Epoch = 1999 Dec. 8.0 TT

T = 1999 Dec. 9.2798 TT  ω = 149.^{o}3039
e = 0.751095  Ω = 246.1346 }2000.0
q = 0.748938 AU  i = 12.8112

a = 3.008934 AU n^{o} = 0.1888362 P = 5.219 years

1999 August 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden